# Kinetic Interacting Particle Langevin Monte Carlo

### v2

@@ 3.2.1 Exponential Integrator (KIPLMC1)
3:for $n=0:K-1$ do
4:     Draw $\varepsilon_{n},\varepsilon^{\prime}_{n}$ from $d_{\theta}+Nd_{x}$-dimensional Gaussians with covariance $C$ given in (9)
5:     $\theta_{n+1}=\theta_{n}+\psi_{1}(\eta)V^{\theta}_{n}-\frac{\psi_{2}(\eta)}{N}\sum_{i=1}^{N}\nabla_{\theta}U(\theta_{n},X^{i}_{n})+\sqrt{\frac{2\gamma}{N}}\varepsilon^{0,\prime}_{n}$
6:     $X^{i}_{n+1}=X^{i}_{n}+\psi_{1}(\eta)V^{x_{i}}_{n}-\psi_{2}(\eta)\nabla_{x}U(\theta_{n},X^{i}_{n})+\sqrt{2\gamma}\varepsilon_{n}^{i,\prime}$
7:     $V^{\theta}_{n+1}=\psi_{0}(\eta)V^{\theta}_{n}-\frac{\psi_{1}(\eta)}{N}\sum_{i=1}^{N}\nabla_{\theta}U(\theta_{n},X^{i}_{n})+\sqrt{\frac{2\gamma}{N}}\varepsilon^{0}_{n}$
8:     $V^{x_{i}}_{n+1}=\psi_{0}(\eta)V^{x_{i}}_{n}-\psi_{1}(\eta)\nabla_{x}U(\theta_{n},X^{i}_{n})+\sqrt{2\gamma}\varepsilon^{i}_{n}$
9:     for $i\in[N]$ return $\theta_{K}$
3.2.2 A Splitting Scheme (KIPLMC2)
The KIPLMC splitting algorithm (KIPLMC2) is an adaptation of the underdamped Langevin sampler introduced by [53], based on classic splitting techniques from MCMC. As with the exponential integrator scheme proposed by [25], the scheme proposed by [53] achieves $\varepsilon$ error in Wasserstein-2 distance using $\varepsilon$ in $\widetilde{\mathcal{O}}(\sqrt{d}/{\varepsilon})$ steps, see, e.g., [53, Table 2]. Specifically, [53] propose an OBABO scheme as the discretisation scheme, which is a second order scheme only requiring the computation of first order derivatives of $U$. The idea is to split the numerical scheme into individually tractable components. In our underdamped case (KIPLD), we have: (A) the update of $\theta$ and $X^{i}$ with known $V^{\theta}$ and $V^{x_{i}}$; (B) update $V^{\theta}$ and $V^{x_{i}}$ with the first order Taylor scheme for the integrals of $\frac{1}{N}\sum_{i=1}^{N}\nabla_{\theta}U(\theta,X^{i})$ and $\nabla_{x_{i}}U(\theta,X^{i})$; (O) solve the Ornstein-Uhlenbeck equation with the terms $V^{\theta}$ and $V^{x_{i}}$. Thus, the algorithm “splits” the equation into components with known solutions. The order in which these steps are taken is critical and in our case we will limit ourselves to considering the OBABO scheme from [53].
In our case we introduce the KIPLMC2 algorithm based on the OBABO scheme, given as
$\theta_{n+1}$ $=\theta_{n}+\eta(\delta V_{n}^{\theta}+\frac{1}{\sqrt{N}}\sqrt{1-\delta^{2}}\varepsilon_{n}^{0})-\frac{\eta^{2}}{2N}\sum_{i=1}^{N}\nabla_{\theta}U(\theta_{n},X_{n}^{i})$
$X^{i}_{n+1}$ $=X^{i}_{n}+\eta(\delta V_{n}^{x_{i}}+\sqrt{1-\delta^{2}}\varepsilon_{n}^{i})-\frac{\eta^{2}}{2}\nabla_{x}U(\theta_{n},X_{n}^{i})$ (KIPLMC2)
$V^{\theta}_{n+1}$ $=\delta^{2}V_{n}^{\theta}-\frac{\delta\eta}{2}\left(\frac{1}{N}\sum_{i=1}^{N}\nabla_{\theta}U(\theta_{n},X_{n}^{i})+\frac{1}{N}\sum_{i=1}^{N}\nabla U(\theta_{n+1},X^{i}_{n+1})\right)+\frac{1}{\sqrt{N}}\sqrt{1-\delta^{2}}(\delta\varepsilon^{0}_{n}-\varepsilon_{n}^{0,\prime})$
$V^{x_{i}}_{n+1}$ $=\delta^{2}V_{n}^{x_{i}}-\frac{\delta\eta}{2}(\nabla_{x}U(\theta_{n},X^{i}_{n})+\nabla_{x}U(\theta_{n+1},X^{i}_{n+1}))+\sqrt{1-\delta^{2}}(\delta\varepsilon^{i}_{n}+\varepsilon^{i,\prime}_{n})$
for $i\in[N]$, where $\delta=e^{-\eta\gamma/2}$ and in this case $\varepsilon^{i}_{n}$ and $\varepsilon^{i,\prime}_{n}$ are i.i.d. standard Gaussians for $i\in[N]$ in this case. The full algorithm is given in Algorithm 2, where we can see the OBABO decomposition more clearly. To produce a half-update of $V^{\theta}$ and $V^{x_{i}}$ the (O) and (B) steps are computed. These are then used to update the parameter and particle estimates (A), followed by the (B) and (O) steps performed in reverse order to update the momenta estimates. Giving us the desirable property of being a second-order, explicit scheme, which does not require computation of $\nabla^{2}U$.
Algorithm 2 KIPLMC2 Algorithm
1:$\gamma,\eta>0$, $N,K\in{\mathbb{N}}$
2:Draw $\theta_{0},X^{i}_{0},V^{\theta}_{0},V^{x_{i}}_{0}$, for $i\in[N]$
3:for $n=0:K-1$ do
4:     Draw $\varepsilon_{n},\varepsilon^{\prime}_{n}$ from $d_{\theta}+Nd_{x}$-dimensional Gaussians
5:     (O+B):
6:     $V^{\theta}_{n+\frac{1}{2}}=\delta V^{\theta}_{n}+\frac{1}{\sqrt{N}}\sqrt{1-\delta^{2}}\varepsilon^{0}_{n}-\frac{\eta}{2N}\sum_{i=1}^{N}\nabla_{\theta}U(\theta_{n},X_{n}^{i})$
7:     $V^{x_{i}}_{n+\frac{1}{2}}=\delta V^{x_{i}}_{n}+\sqrt{1-\delta^{2}}\varepsilon^{i}_{n}-\eta\nabla_{x}U(\theta_{n},X_{n}^{i})$
8:     (A):
9:     $\theta_{n+1}=\theta_{n}+\eta V^{\theta}_{n+\frac{1}{2}}$
10:     $X^{i}_{n+1}=X^{i}_{n}+\eta V^{x_{i}}_{n+\frac{1}{2}}$
11:     (B+O):
12:     $V^{\theta}_{n+1}=\delta(V^{\theta}_{n+\frac{1}{2}}-\frac{\eta}{2N}\sum_{i=1}^{N}\nabla_{\theta}U(\theta_{n+1},X_{n+1}^{i})+\frac{1}{\sqrt{N}}\sqrt{1-\delta^{2}}\varepsilon^{0,\prime}_{n}$
13:     $V^{x_{i}}_{n+1}=\delta(V^{x_{i}}_{n+\frac{1}{2}}-\frac{\eta^{2}}{2}\nabla_{x}U(\theta_{n+1},X^{i}_{n+1})+\frac{1}{\sqrt{N}}\sqrt{1-\delta^{2}}\varepsilon^{0,\prime}_{n}$ return $\theta_{K}$
@@ 4 Nonasymptotic Analysis
In this section, we provide the convergence results for both the analytic and numerical schemes in the nonasymptotic regime. In Section 4.1, we lay out our assumptions about our target measure. In Section 4.2, we introduce the proof strategy for reader’s convenience, which can be also useful for extending and proving similar results leveraging the same structure as we do. In Section 4.3, we provide the full nonasymptotic bounds – in particular, in Section 4.3.1, we identify the stationary measure for KIPLD and show an exponential convergence rate to it, as well as, its concentration onto the MMLE solution $\bar{\theta}_{\star}$ as $N$ grows. Following this, in Section 4.3.2, the error bounds for the two algorithms are provided, thus allowing us to identify the convergence rate of KIPLMC1 and KIPLMC2 to the MMLE solution $\bar{\theta}_{\star}$, which is discussed in 4.4.
@@ 4.1 Assumptions
We first lay out our assumptions to prove the convergence of the numerical schemes outlined in 3. Our assumptions are generic for the analysis of Langevin diffusions, i.e., we assume strong convexity and $L$-Lipschitz gradients for the potential $U$. These assumptions are akin to the assumptions made in [31, 23, 24] for proving the convergence of ULA. As such, the assumptions are necessary for basic building blocks of the theory – but it is possible to relax them, as can be seen in [71, 1].
We remark nonetheless that, while in the works cited above, assumptions correspond to strong log-concavity (and Lipschitz regularity) of the target measure (i.e. strongly convex $U$ implies that the target measure is assumed to be strongly log-concave), in our setting this corresponds to imposing a strong structure on the joint statistical model $p_{\theta}(x,y)$ in $\theta$ and $x$. Note that, in our case the potential is given by $U(\theta,x)=-\log p_{\theta}(x,y)$, in other words, our potential is not merely the log-target, it is the joint negative log-likelihood. Thus, the assumptions below require statisticians to check whether their models satisfy these assumptions. We check in the experimental section that for some generic models such as Bayesian logistic regression, this is possible. As noted above, however, these strong assumptions on the statistical model should not be difficult to relax using the standard non-log-concave sampling results – but this direction is out of scope of the present work which considers the strongly convex case.
@@
Let $z,z^{\prime}\in\mathbb{R}^{d_{\theta}+d_{x}}$, we suppose that there exists a $\mu$ s.t.
@@
This assumption also implies that $\nabla^{2}U$ is positive definite, i.e., $\nabla^{2}U(z)\succeq\mu I$ for all $z\in\mathbb{R}^{d_{\theta}+d_{x}}$.
This assumption is equivalent to an assumption of joint strong convexity in $\theta$ and $x$ for $U$, which ensures exponentially fast convergence of the KIPLD to the global minimiser and places a quadratic lower growth bound on $U$. This assumption also means that we can apply the Leibniz differentiation rule under integration for the marginal of the probability model $p_{\theta}(x,y)=e^{-U(\theta,x)}$ [8, Theorem 16.8]. On a more statistical note, this assumption requires that the log-likelihood function $\log p_{\theta}(x,y)$ should be strongly concave in both the latent variable $x$ and the parameter $\theta$.
Next, we provide the regularity assumption on $U$.
@@
We suppose $U\in\mathcal{C}^{1}$ and for any $z,z^{\prime}\in\mathbb{R}^{d_{\theta}+d_{x}}$ there exists a constant $L>0$ s.t.
@@
This assumption is a common assumption in the analysis of stochastic differential equations, as it ensures stability of the diffusion KIPLD, leading to strong solutions, and stable numerical discretisations (see e.g. [55]). Note also that from the perspective of the probabilistic model, this assumption is the same as assuming that $\log p_{\theta}(x,y)$ is gradient Lipschitz. Note that, it is possible to also relax these assumptions, incorporating superlinearly growing gradients, see, e.g., [45].
@@ 4.2 The proof strategy
Given the assumptions above, we aim at bounding the optimisation error of the numerical schemes, i.e., the difference between the law of the numerical scheme and the optimal MMLE solution $\bar{\theta}_{\star}$. Let $(\theta_{n})_{n\geq 0}$ be the sequence of iterates generated by a numerical scheme, for example, KIPLMC1 or KIPLMC2. Finally, let $\pi_{\Theta}$ denote the $\theta$-marginal of the stationary measure of the diffusion KIPLD, and $\delta_{\bar{\theta}_{\star}}$ denote the Dirac measure at $\bar{\theta}_{\star}$. We denote the law of $\theta_{n}$ by $\mathcal{L}(\theta_{n})$. The optimisation error is then defined as ${\mathbb{E}}[\|\theta_{n}-\bar{\theta}_{\star}\|^{2}]^{1/2}$.
In order to proceed, we first note that ${\mathbb{E}}[\|\theta_{n}-\bar{\theta}_{\star}\|^{2}]^{1/2}=W_{2}(\mathcal{L}(\theta_{n}),\delta_{\bar{\theta}_{\star}})$, where $\mathcal{L}(\theta_{n})$ denotes the law of $\theta_{n}$. This is due to the fact that the set of couplings between a measure $\nu$ and another measure $\delta_{y}$ contains a single element $\delta_{y}\otimes\nu$ [61, Section 1.4], thus the infimum in Wasserstein-2 distance is attained by the coupling $\delta_{\bar{\theta}_{\star}}\otimes\mathcal{L}(\theta_{n})$. Using this, we can then write
$\leq\underbrace{W_{2}(\pi_{\Theta},\delta_{\bar{\theta}_{\star}})}_{\textnormal{concentration}}+\underbrace{W_{2}(\mathcal{L}(\theta_{n}),\pi_{\Theta})}_{\textnormal{convergence}},$ (10)
using the triangle inequality as the Wasserstein distance is a metric. The first term in the right-hand side of (10) is the concentration of the stationary measure $\pi_{\Theta}$ onto the MMLE solution $\bar{\theta}_{\star}$, which is given by Proposition 2 below. The second term is the convergence of the numerical scheme to the stationary measure $\pi_{\Theta}$ which will be proved for each scheme separately.
4.3 Nonasymptotic convergence bounds
As outlined above in (10), we will first show the concentration of the stationary measure $\pi_{\Theta}$ onto the MMLE solution $\bar{\theta}_{\star}$. We then show the convergence of the numerical schemes to the stationary measure $\pi_{\Theta}$. Finally, we will provide the error bounds for the numerical schemes.

### v3

@@ 3.2 A Splitting Scheme (KIPLMC2)
3:for n=0:M1n=0:M-1 do
4:  Draw $\varepsilon_{n},\varepsilon^{\prime}_{n}$ from $d_{\theta}+Nd_{x}$-dimensional Gaussians and for $i\in[N],$
5:  Compute the gradients of the potential:
$F_{n}^{\theta}$ $=\frac{1}{N}\sum_{i=1}^{N}\nabla_{\theta}U(\theta_{n},X_{n}^{i}),\qquad F_{n}^{x_{i}}=\nabla_{x}U(\theta_{n},X_{n}^{i}).$
6:  Update the momenta via (O)+(B) steps:
$V_{n+\frac{1}{2}}^{\theta}$ $=\delta V_{n}^{\theta}+\sqrt{\frac{1-\delta^{2}}{N}}\xi_{n}^{0}-\frac{\eta}{2}F_{n}^{\theta},$
$V_{n+\frac{1}{2}}^{x_{i}}$ $=\delta V_{n}^{x_{i}}+\sqrt{1-\delta^{2}}\xi_{n}^{i}-\frac{\eta}{2}F_{n}^{x_{i}},$
7:  Update parameters and particles via (A):
$\theta_{n+1}$ $=\theta_{n}+\eta V_{n+\frac{1}{2}}^{\theta},\qquad X_{n+1}^{i}=X_{n}^{i}+\eta V_{n+\frac{1}{2}}^{x_{i}},$
for $i\in[N]$.
8:  Re-compute the gradient potentials
$F_{n+1}^{\theta}=\frac{1}{N}\sum_{i=1}^{N}\nabla_{\theta}U(\theta_{n+1},X_{n+1}^{i}),\qquad F_{n+1}^{x_{i}}=\nabla_{x}U(\theta_{n+1},X_{n+1}^{i}).$
9:  Update the momenta via the (B)+(O) steps:
$V_{n+1}^{\theta}$ $=\delta V_{n+\frac{1}{2}}^{\theta}-\frac{\delta\eta}{2}F_{n+1}^{\theta}+\sqrt{\frac{1-\delta^{2}}{N}}\xi_{n}^{0,\prime},$
$V_{n+1}^{x_{i}}$ $=\delta V_{n+\frac{1}{2}}^{x_{i}}-\frac{\delta\eta}{2}F_{n+1}^{x_{i}}+\sqrt{1-\delta^{2}}\xi_{n}^{i,\prime},$
for $i\in[N]$. return $\theta_{M}$
In the OBABO case, over a full time-step $\eta$, the (O) step is applied over a $\eta/2$ half-step, the (B) step is also applied to a $\eta/2$ half-step, (A) is applied as a full time-step $\eta$, followed by a reversed application of (B) and (O) for half-steps. We denote these intermediate time-steps by $n+\frac{1}{4}$, $n+\frac{1}{2}$ and $n+\frac{3}{4}$, though we note that these time-steps do not correspond to simulated time. The order in which these steps are taken is critical [41] and in our case we will limit ourselves to considering the OBABO ordering.
More precisely, the Ornstein-Uhlenbeck half-step (O) flow is given as
$\mathrm{d}{\mathbf{V}}_{t}^{\theta}$ $=-\gamma{\mathbf{V}}_{t}^{\theta}\mathrm{d}t+\sqrt{\frac{2\gamma}{N}}\mathrm{d}{\mathbf{B}}_{t}^{0},$
$\mathrm{d}{\mathbf{V}}_{t}^{x_{i}}$ $=-\gamma{\mathbf{V}}_{t}^{x_{i}}\mathrm{d}t+\sqrt{2\gamma}\mathrm{d}{\mathbf{B}}_{t}^{i},$
over the time-interval $t_{n},t_{n+1/2}]$, for $i\in[N]$, where the parameters $\bm{\theta}_{t}$ and ${\mathbf{X}}_{t}^{i}$ are frozen and we initialise at $({\mathbf{V}}_{n\eta}^{\theta},{\mathbf{V}}_{n\eta}^{x_{i}})=(V_{n}^{\theta},V_{n}^{x_{i}})$. Using the classic analytic solution to the Ornstein-Uhlenbeck equation and the Itô Isometry we obtain,
$V_{n+\frac{1}{4}}^{\theta}$ $=\delta V_{n}^{\theta}+\frac{1}{\sqrt{N}}\sqrt{1-\delta^{2}}\xi_{n}^{0},$ (O)
$V_{n+\frac{1}{4}}^{x_{i}}$ $=\delta V_{n}^{x_{i}}+\sqrt{1-\delta^{2}}\xi_{n}^{i},$
with $\delta=e^{-\eta\gamma/2}$ and i.i.d. standard Gaussians $\xi_{n}^{0}$ and $\xi_{n}^{i}$, in $\mathbb{R}^{d_{\theta}}$ and $\mathbb{R}^{d_{x}}$, respectively.
The momentum drift update (B) updates $V^{\theta}$ and $V^{x_{i}}$ with the first order Taylor scheme for the integrals of $N^{-1}\sum_{i=1}^{N}\nabla_{\theta}U(\theta,X^{i})$ and $\nabla_{x_{i}}U(\theta,X^{i})$, i.e. we consider the flow to be given as
$\mathrm{d}{\mathbf{V}}_{t}^{\theta}$ $=-F_{n}^{\theta}\mathrm{d}t,$
$\mathrm{d}{\mathbf{V}}_{t}^{x_{i}}$ $=-F_{n}^{x_{i}}\mathrm{d}t,$
over $t\in[t_{n},t_{n+1/2}]$, where we recall the definitions of $F_{n}^{\theta}$ and $F_{n}^{x_{i}}$ from Sec. 3.1 and we initialise at $({\mathbf{V}}_{n\eta}^{\theta},{\mathbf{V}}_{n\eta}^{x_{i}})=(V_{n+\frac{1}{2}}^{\theta},V_{n+\frac{1}{2}}^{x_{i}})$. This results in the update,
$V_{n+\frac{1}{2}}^{\theta}$ $=V_{n+\frac{1}{4}}^{\theta}+\frac{\eta}{2}F_{n}^{\theta},$ (B)
$V_{n+\frac{1}{2}}^{x_{i}}$ $=V_{n+\frac{1}{4}}^{x_{i}}+\frac{\eta}{2}\nabla_{x}F_{n}^{x_{i}}.$
Finally, the (A) update considers the partial flow,
$\mathrm{d}\bm{\theta}_{t}$ $={\mathbf{V}}_{t}^{\theta}\mathrm{d}t,$
$\mathrm{d}{\mathbf{X}}_{t}^{i}$ $={\mathbf{V}}_{t}^{x_{i}}\mathrm{d}t,$
over $t\in[t_{n},t_{n+1}]$, where we consider $({\mathbf{V}}_{t}^{\theta},{\mathbf{V}}_{t}^{x_{i}})$ is frozen at $(V_{n+1/2}^{\theta},V_{n+1/2}^{x_{i}})$ and we initialise at $(\bm{\theta}_{n\eta},{\mathbf{X}}_{n\eta}^{i})=(\theta_{n},X_{n}^{i})$. Hence, the (A) update is given as,
$\theta_{n+1}$ $=\theta_{n}+\eta V_{n+\frac{1}{2}}^{\theta},$ (A)
$X_{n+1}^{i}$ $=X_{n}^{i}+\eta V_{n+\frac{1}{2}}^{x_{i}}.$
The algorithm is then completed by repeating steps (B) and (O),
$V_{n+1}^{\theta}$ $=\delta\left(V_{n+\frac{1}{2}}^{\theta}-\frac{\eta}{2}F_{n+1}^{\theta}\right)+\sqrt{\frac{1-\delta^{2}}{N}}\xi_{n}^{0,\prime},$ (B+O)
$V_{n+1}^{x_{i}}$ $=\delta\left(V_{n+\frac{1}{2}}^{x_{i}}-\frac{\eta}{2}F_{n+1}^{x_{i}}\right)+\sqrt{1-\delta^{2}}\xi_{n}^{i,\prime},$
over the interval $t\in[t_{n+1/2},t_{n+1}]$ for standard Gaussians $\xi_{n}^{0,\prime}$ and $\xi_{n}^{i,\prime}$ in $\mathbb{R}^{d_{\theta}}$ and $\mathbb{R}^{d_{x}}$ respectively.
The full updates are given in Algorithm 2 and a full derivation of the KIPLMC2 algorithm is given in Appendix B.
@@ 4 Nonasymptotic Analysis
In this section, we provide the convergence results for both the analytic and numerical schemes in the nonasymptotic regime. In Section 4.1, we lay out our assumptions about our target measure. In Section 4.2, we introduce the proof strategy for the reader’s convenience, which may be useful for extending and proving similar results. In Section 4.3, we provide the full nonasymptotic bounds – in particular, in Section 4.3.1, we identify the stationary measure for KIPLD and show an exponential convergence rate to it, as well as, its concentration onto the MMLE solution $\bar{\theta}_{\star}$ as $N$ grows. Following this, in Section 4.3.2, the error bounds for the two algorithms are provided, allowing us to identify the convergence rate of KIPLMC1 and KIPLMC2 to the MMLE solution $\bar{\theta}_{\star}$.
@@ 4.1 Assumptions
Our assumptions are generic for the analysis of Langevin diffusions, i.e., we assume strong convexity and $L$-Lipschitz gradients for the potential $U$. These assumptions are akin to the assumptions made in Durmus and Moulines [24], Dalalyan [20], Dalalyan and Karagulyan [18] for proving the convergence of ULA. It is, however, possible to relax them, as can be seen in Zhang et al. [53], Akyildiz and Sabanis [3].
@@
Suppose $U\in C^{2}$. Let $z,z^{\prime}\in\mathbb{R}^{d_{\theta}+d_{x}}$, we suppose that there exists a $\mu$ s.t.
@@
From this follows that $\nabla^{2}U$ is positive definite, i.e., $\nabla^{2}U(z)\succeq\mu I$ for all $z\in\mathbb{R}^{d_{\theta}+d_{x}}$.
This assumption is equivalent to an assumption of joint strong convexity in $\theta$ and $x$ for $U$, which ensures exponentially fast convergence of the KIPLD to the global minimiser and places a quadratic lower growth bound on $U$. This assumption also means that we can apply the Leibniz differentiation rule under integration for the marginal of the probability model $p_{\theta}(x,y)=e^{-U(\theta,x)}$ [8, Theorem 16.8]. Next, we provide a smoothness assumption on the potential $U$.
@@
We suppose for any $z,z^{\prime}\in\mathbb{R}^{d_{\theta}+d_{x}}$ there exists a constant $L>0$ s.t.
@@
This assumption is a common assumption in the analysis of stochastic differential equations, as it ensures stability of the diffusion KIPLD, leading to strong solutions, and stable numerical discretisations [34]. Observe that this assumption places an upper quadratic growth bound on $U$ or, from the perspective of the probabilistic model, $\log p_{\theta}(x,y)$ is of quadratic growth in both $\theta$ and $x$.
We remark nonetheless that, in the works cited above, the assumptions are on the strong log-concavity of the target measure. In our setting, this corresponds to imposing a structure on the joint statistical model $p_{\theta}(x,y)$ in $\theta$ and $x$, as our potential is given as $U(\theta,x)=-\log p_{\theta}(x,y)$. This requires some care on the user side as a statistical model with strongly log-concave posterior $p_{\theta}(x|y)$ (with fixed $\theta$) may not necessarily satisfy strong log-concavity of $p_{\theta}(x,y)$ in $(x,\theta)$ jointly. We verify in the experimental section that for some generic models such as Bayesian logistic regression, this is possible. As noted above, however, these strong assumptions on the statistical model should not be difficult to relax using the standard non-log-concave sampling results – but this direction is out of scope of the present work which considers the strongly convex case.
@@ 4.2 The proof strategy
Let $(\theta_{n})_{n\geq 0}$ be the sequence of iterates generated by a numerical scheme, for example, KIPLMC1 or KIPLMC2. Let $\pi_{\Theta}$ denote the $\theta$-marginal of the stationary measure of the diffusion KIPLD, and $\delta_{\bar{\theta}_{\star}}$ denote the Dirac measure at $\bar{\theta}_{\star}$. Finally, we denote the law of $\theta_{n}$ by $\mathcal{L}(\theta_{n})$.
In order to proceed, we first note that the optimisation error is given as ${\mathbb{E}}[\|\theta_{n}-\bar{\theta}_{\star}\|^{2}]^{1/2}=W_{2}(\mathcal{L}(\theta_{n}),\delta_{\bar{\theta}_{\star}})$. This is due to the fact that the set of couplings between a measure $\nu$ and another measure $\delta_{y}$ contains a single element $\delta_{y}\otimes\nu$ [45, Section 1.4], thus the infimum in Wasserstein-2 distance is attained by the coupling $\delta_{\bar{\theta}_{\star}}\otimes\mathcal{L}(\theta_{n})$. This implies that ${\mathbb{E}}[\|\theta_{n}-\bar{\theta}_{\star}\|^{2}]^{1/2}=W_{2}(\mathcal{L}(\theta_{n}),\delta_{\bar{\theta}_{\star}})$ and consequently using triangle inequality
${\mathbb{E}}[\|\theta_{n}-\bar{\theta}_{\star}\|^{2}]^{1/2}\leq\underbrace{W_{2}(\mathcal{L}(\theta_{n}),\pi_{\Theta})}_{\textnormal{convergence}}+\underbrace{W_{2}(\pi_{\Theta},\delta_{\bar{\theta}_{\star}})}_{\textnormal{concentration}},$ (8)
as the Wasserstein distance is a metric. The first term is the convergence of the numerical scheme to the stationary measure $\pi_{\Theta}$ which will be proved for each scheme separately. The second term in the right-hand side of (8) is the concentration of the stationary measure $\pi_{\Theta}$ onto the MMLE solution $\bar{\theta}_{\star}$, which is given by Proposition 2 below.
4.3 Convergence bounds
As outlined above in (8), we will first show the concentration of the stationary measure $\pi_{\Theta}$ onto the MMLE solution $\bar{\theta}_{\star}$, then show the convergence of the KIPLD to the stationary measure $\pi_{\Theta}$. Finally, we will provide the error bounds for the numerical schemes.